OpenStudy (anonymous):

students in an english class took a final exam. they took equivalent forms of the exam at monthly intervals thereafter. the average score S(t), in percent , after t months was foundto be given by: S(t)=68-20log(t+1), t is greater than 0. 1a). what was the average score when they initially took the test, t=0 1b) what was the average score after 14 months 1c) after what time was the average score 40%

OpenStudy (anonymous):

since S(t) is the average score for t months a) S(0) : plug in 0 for every t in S(t). Should give S(t) = 68 b) S(14): plug in 14 for every t in S(t) [only one t in this case] c) S(t) = 0.4. Find t. lmk if u need help with this one after doing the 2 first

OpenStudy (anonymous):

i dont really understand

OpenStudy (anonymous):

which one?

OpenStudy (anonymous):

heres what i have 1a) 68

OpenStudy (anonymous):

the other ones are confusing

OpenStudy (anonymous):

b). plug in 14 for every t S(14) = 68 -20log(14+1) = 68 - 20(1.176) = 44.478 the value of log15 was obtained with my calculator. note: i used log base 10 [log button] and not the natural logarithm [ln button]

OpenStudy (anonymous):

and 1c.

OpenStudy (anonymous):

0.4 = 68 - 20log(t+1) re-arranged 20log(t+1) = 67.6 log(t+1) = 3.38 assuming log is log base 10 and not log base e, doing it's inverse operation: (t+1) = 10^(3.38) t = 2397.83 which seems like a very high number.

OpenStudy (anonymous):

thank you.